OpenStudy (anonymous):

limit as x approaches 0 from the left of : e^(1/x) when x!=0. c when x= 0. Find the value of c that will make function continuous

OpenStudy (anonymous):

Do I find the limit as e^(1/x) approaches 0 from the left and let that eqaul c?

OpenStudy (anonymous):

\[f(x)=\begin{cases} e^{1/x}&\text{for }x\not=0\\ c&\text{for }x=0 \end{cases}\] And you're asked to find \(c\) that makes \(f(x)\) continuous, i.e. such that \(\displaystyle\lim_{x\to0^-}f(x)=c.\) Is that right?

OpenStudy (anonymous):

Wait, I have that wrong...

OpenStudy (anonymous):

There is no value of \(c\) that makes the function continuous, since \(\displaystyle\lim_{x\to0}f(x)\) doesn't exist. Also, \(\displaystyle\lim_{x\to0^-}f(x)=0\).